# Sensing Gravity with Polarized Electromagnetic Radiation

### v1

@@ VIII.5 Using Polarization Wiggling to Measure the State Parameters of a Plane Gravitational Wave
The effect of the other 6 parameters are entangled in the amplitude $W$ and phase $\gamma$. Assuming the polarized radiation is a plane electromagnetic wave, both the wiggling amplitude $W$ and phase $\gamma$ are measurable quantities. It should therefore be sufficient, at least in principle, to determine the remaining 6 parameters of the state vector $\mathcal{K}$ by measuring the wiggling amplitudes and phases of three beams of polarized radiation emitted by emitters at known positions from different directions.
@@ IX Conclusions
Spacetimes with conformally flat backgrounds is an extensive class of spacetimes that includes most spacetimes of interest in cosmology. They have one thing in common: their homogeneous backgrounds are conformally related to 4-dimensional Minkowski space. This relationship  gives rise to  a major simplification in the analysis of polarization wiggling: We showed how it is possible to calculate the polarization wiggle rates in the conformal frame and transform the results back to the physical frame. The conformal frame is a much simpler geometry than the physical frame due to its flat Minkowski background, and the wiggle rates in the physical frame can be obtained by simply multiplying the conformal frame results with a scale factor. This simplification arises when electromagnetic radiation is decomposed into plane waves. As electromagnetic plane waves propagate along null geodesics, we can make use of the fact that null geodesics are invariant under conformal transformations and that Maxwell’s equations in a 4-dimensional vacuum are also conformally invariant.
In linear gravity, the gravitational field, represented by the metric perturbation $h_{\mu\nu}$, can be decomposed into scalar, vector and tensor perturbations that are sourced and evolve independently. The polarization wiggle rates induced by scalar, vector and tensor perturbations to the metric are gauge invariant quantities that can be analyzed independently.
@@
Vector perturbations to the metric, on the other hand, may induce polarization wiggling. The polarization axis wiggle rate $\omega^{(V)}$ is a direct measure of the longitudinal gravitomagnetic field $\mathcal{B}$ and thereby the matter currents sourcing it. The contribution to the polarization axis wiggle rate from vector perturbations has two terms. The first contribution is due to the difference in the angular frame dragging rate around the direction of propagation between the frame of emission and the frame of observation. The second contribution comes from the presence of non-stationary gravitomagnetic fields along the radiation trajectory. This implies that, for a stationary gravitomagnetic field, the polarization axis wiggle rate simply measures the difference in angular frame dragging rate between emission and observation. Thus, stationary gravitomagnetic fields will not induce polarization wiggling of any passing electromagnetic radiation if this radiation was emitted and measured at locations with negligible gravitomagnetic fields.
@@
Tensor perturbations to the metric also induce polarization wiggling. The effect of a single gravitational plane wave (tensor mode) with an arbitrary state of polarization was studied in detail for two cases: A tensor mode propagating on respectively a flat background and in an expanding, matter-dominated universe. In both cases, we showed how the state parameters of the tensor mode can be determined by measurements of the polarization wiggling frequency, amplitude and phase of radiation emitted by emitters at known positions from at least three different directions. In both cases, the polarization axis of the observed radiation will wiggle with a frequency that equals the frequency of the gravitational wave. Both amplitude and phase of the polarization wiggling can be used to determine the properties of the tensor mode. The amplitude and phase of the polarization wiggling vary with the position of the emitter; the polarization wiggle rate is sensitive to both direction of the electromagnetic radiation used for the measurement and the distance to the emitter. The wiggling amplitude vanishes for gravitational waves with wave lengths much larger than the distance to the emitter.

### v2

@@ VIII.5 Using Polarization Wiggling to Measure the State Parameters of a Plane Gravitational Wave
The effect of the other 6 parameters are entangled in the amplitude $W$ and phase $\gamma$. Assuming the polarized radiation is a plane electromagnetic wave, both the wiggling amplitude $W$ and phase $\gamma$ are measurable quantities. It should therefore, in principle, be sufficient to determine the remaining 6 parameters of the state vector $\mathcal{K}$ by measuring the wiggling amplitudes and phases of three beams of polarized radiation emitted by emitters at known positions from different directions.
@@ IX Conclusions
Spacetimes with conformally flat backgrounds is an extensive class of spacetimes that includes most spacetimes of interest in cosmology. They have one thing in common: their homogeneous backgrounds are conformally related to 4-dimensional Minkowski space. This relationship  gives rise to  a major simplification in the analysis of polarization wiggling: We showed how it is possible to calculate the polarization wiggle rates in the conformal frame and transform the results back to the physical frame. The conformal frame is a much simpler geometry than the physical frame due to its flat Minkowski background, and the wiggle rates in the physical frame can be obtained by just multiplying the conformal frame results with a scale factor. This simplification arises when electromagnetic radiation is decomposed into plane waves. Within the geometric optics limit, electromagnetic plane waves propagate along null geodesics. Therefore, we can make use of the fact that null geodesics are invariant under conformal transformations and that Maxwell’s equations in a 4-dimensional vacuum are also conformally invariant.
In linear gravity, the gravitational field, represented by the metric perturbation $h_{\mu\nu}$, can be decomposed into scalar, vector and tensor perturbations that are sourced and evolve independently. We found that the polarization wiggle rates induced by scalar, vector and tensor perturbations to the metric are gauge invariant quantities that can be analyzed independently.
@@
Vector perturbations to the metric, on the other hand, may induce polarization wiggling. The vector-induced polarization axis wiggle rate $\omega^{(V)}$ is a direct measure of the longitudinal gravitomagnetic field $\mathcal{B}$ and thereby the matter currents sourcing it. The contribution to the polarization axis wiggle rate from vector perturbations has two terms. The first term is the difference in the angular frame dragging rate around the direction of propagation between the frame of emission and the frame of observation. The second term integrates the angular acceleration of the frame dragging rate caused by non-stationary gravitomagnetic fields present along the radiation trajectory. This implies that, for a stationary gravitomagnetic field, the polarization axis wiggle rate simply measures the difference in angular frame dragging rate between emission and observation. Thus, stationary gravitomagnetic fields will not induce polarization wiggling of any passing electromagnetic radiation if this radiation was emitted and measured at locations with negligible gravitomagnetic fields.
@@
Tensor perturbations to the metric also induce polarization wiggling. We studied the effect of a single gravitational plane wave (tensor mode) with an arbitrary state of polarization. Two particular cases were investigated in detail: A tensor mode propagating on a flat background and in an expanding, matter-dominated universe. In both cases, we showed how the state parameters of the tensor mode can be determined by measurements of the polarization wiggling frequency, amplitude and phase of electromagnetic radiation emitted by emitters at known positions from at least three different directions. In both cases, the polarization axis of the observed electromagnetic radiation will wiggle with a frequency that equals the frequency of the gravitational wave. Both amplitude and phase of the polarization wiggling can be used to determine the properties of the tensor mode. The amplitude and phase of the polarization wiggling vary with the position of the emitter; the polarization wiggle rate is sensitive to both direction of the electromagnetic radiation and distance to the emitter. The wiggling amplitude vanishes for gravitational waves with wave lengths much larger than the distance to the emitter.